# Burlingame, CA 94010, USA

### v2

@@ Abstract
Apache Spark is better suited for load data mining and machine learning that require a lot of iteration by using memory-distributed data sets. Due to the complexity of Spark, the high learning threshold of Scala, and the low reusability of its code, this paper designs and implements a Spark-based visual process AI+machine learning method under a big data environment. On the one hand, it designs component models to describe the basic steps of machine learning, including data preprocessing, feature processing, and model training. Practice and validate evaluation. On the other hand, a visual process modeling tool is provided to support analysts to design machine learning processes, which can be translated automatically into Spark platform code for efficient execution. This tool can greatly improve the AI machine learning efficiency of the Spark platform. This paper introduces the method theory, key technologies, and effectiveness of the tool.
This paper explores the application of Spark in the field of large model agents. Langchain, as an open-source framework, is committed to simplifying the development of end-to-end applications based on language models. It provides interfaces for interacting with a variety of large language models, optimizing prompt engineering, and endowing large models with the ability to invoke external tools. LangGraph demonstrates its powerful state management and graph construction capabilities by defining node functions and graphs to build complex agent applications. The development of Spark agent applications based on LangGraph has further promoted the development of AI applications in the big data analysis environment .

### v3

@@ Abstract
This paper introduces a visual process modeling tool for AI and machine learning in big data contexts, utilizing the LangGraph framework to construct intelligent Spark agents. The tool represents key machine learning stages—data preprocessing, feature engineering, model training, and evaluation—as modular components. Analysts can visually design workflows, which are then automatically translated into optimized Spark code for execution. This approach simplifies the complexity of Apache Spark, reduces the learning curve associated with Scala, and enhances code reusability. The paper discusses the theoretical foundations, key technologies, and evaluates the effectiveness of the proposed solution.